# EPIM: Efficient Processing-In-Memory Accelerators based on Epitome

### v2

@@ Abstract
Index Terms:
Processing-In-Memory, Model Compression, Efficient Neural Networks
* Equal contribution.
[…]
I Introduction
The emergence of Processing-In-Memory (PIM) accelerators offers a promising approach to boost the computational efficiency of neural network operations [1]. Despite the advances in current PIM accelerators, they still face certain challenges. One challenge arises from the higher cost of writing to emerging memory compared to reading from it. Due to this limitation, PIM accelerators typically require loading all neural network weights onto memristor crossbars prior to conducting computations. However, modern large-scale neural network models have seen a significant increase in parameter sizes. Even for the existing PIM accelerators with maximum memristors, there still exists a significant challenge in deploying a standard backbone model like ResNet-50[2]. As such, the deployment of advanced Convolutional Neural Networks (CNN) models on PIM is hindered.
To address this challenge, researchers have proposed methods that fall into two main categories. Various methods, including pruning and quantization, aim to represent neural operator weights using a reduced number of parameters[3, 4]. Conversely, approaches such as neural architecture search concentrates on identifying optimal combinations of neural operators[5, 6]. Despite the advancements, the majority of these studies have focused on adapting existing convolution operators. The crucial endeavor of developing compact and high-performance operators specifically tailored for PIM accelerators has largely been overlooked.
@@ I Introduction
To craft such neural operator, we began by reviewing existing literature. Prior research highlights the potential of leveraging overlaps in convolution kernels, introducing a compact neural operator termed “epitome”[7]. This operator reconstructs convolution kernels from a smaller epitome tensor. A more compact neural network can be crafted by replacing the its convolutions with epitomes. We identified that, with software and hardware adjustments, epitome could be adapted to PIM accelerators. Specifically, from the hardware side, epitome only introduces minor modifications to the data path. Hence, we adopted epitome as our foundation. By integrating epitome with our proposed methods, we achieved notable compression rates while maintaining a high accuracy.
@@
II Related work
II-A Process-In-Memory Architecture for CNN Acceleration
Commonly, PIM architecture consists of multiple memristor crossbars. The weight matrices of CNN are mapped as the conductance of the memristor crossbar cells. However, the writing latency of the memristor crossbar cell is multiple times larger than the reading latency[1]. Before inference, the weight matrices of all CNN layers need to be mapped onto memristor crossbars. Such restrictions stimulate the need to support model compression algorithms on PIM architecture.
Figure 1: Epitome’s Reconstruction into Convolution
II-B Epitome
@@
$\small E_{s}=E_{i}[p:p+w,q:q+h,c_{in}:c_{in}+\beta_{1},c_{out}:c_{out}+\beta_{2}],$ (1)
where $(p,q,c_{in},c_{out})$ denote the starting index of the sub-tensor in the epitome and $(w,h,\beta_{1},\beta_{2})$ denotes the length of the sub-tensor along each dimension. The sampling process is repeated until the fetched small pieces of weight tensors can be concatenated to match the dimensions of the convolution. Figure 1 is a visual illustration of such process. Unlike convolutions, whose shapes are determined by the input and output channels, epitomes have greater flexibility in shape. They can be of four-dimensional tensors with any size, as long as their sampled patches can reconstruct the convolutions[7]. Epitome-based neural networks can significantly reduce parameter size compared to traditional CNNs, with only minimal impact on algorithmic performance[7].
II-C Quantization
Quantization methods reduce model size by using low bitwidth for weights and activations in neural networks[8]. In quantization, the quantizer maps weights and activations into integers with a real-valued scaling factor $S$ and an integer zero point $Z$ as follows:
$\small Q(r)=\text{Int}\big{(}{r}/{S}\big{)}-Z$ (2)
@@ II-C Quantization
$\small S=\frac{\beta-\alpha}{2^{k}-1}$ (3)
@@
III Framework Overview
Figure 2: Illustration of EPIM Framework.
@@ III Framework Overview
IV EPIM Design Method
This section introduces our method EPIM. Section IV-A shows the adjustments of epitome on PIM accelerators. Section IV-B presents details of our quantization method. Section IV-C describes the specifications of our PIM accelerator.
IV-A Mapping and Adjusting Epitomes for PIM Accelerators
Just like convolutions, epitomes are four-dimensional tensors that can be mapped onto memristors in a similar manner. The distinction between epitomes and convolutions arises during inference: while all word lines and bit lines corresponding to a convolution’s weight are activated all together, an epitome causes a memristor crossbar to be activated multiple times. Each activation only engages the word lines and bit lines associated with small patches sampled from the epitome. In our implementation, we use the same mapping strategy as [9].
Motivated by the size flexibility of the epitomes, we can adjust their shapes to better utilize memristors. Specifically, we aim for $c_{out}$ and $c_{in}\times p\times q$ to align as integral multiples of the crossbar size. Following the mapping approach from [9], where $c_{in}$, $p$ and $q$ dimensions correspond to word lines, while $c_{out}$ dimension corresponds to bit lines on memristor crossbars. By setting to these dimensions to desired values, we ensure epitomes make full use of memristor crossbars.
IV-B Quantization Method for Epitome
When directly quantizing the epitome with ultra-low precision, we observed a notable decrease in accuracy. Table II indicates that while the 5-bit quantized weights for ResNet-50’s epitome achieve 73.59% accuracy, this drops to 69.95% for 3-bit weights. To enhance accuracy under such low-bit quantization, we suggest two adjustments: Firstly, given the parallel computation between PIM accelerator crossbars, we allocate a scaling factor to each crossbar. Secondly, we observe that the epitome reconstructs convolution by sampling overlapping small patches, with certain patches from the epitome being repeated more often than others in the reconstructed convolution. When quantizing the epitome, it’s logical to assign different importance levels to different parts based on their repetition frequency. As depicted in Figure 2(c), we notice that the center parts (green) of epitomes are usually repeated more frequently than the other parts (blue). We calculate the weighted sum of the epitome’s maximum and minimum values in the center parts and other parts as quantization range $\alpha$ and $\beta$, based on these repetitions. The detailed calculation process is as follows:
$\small\alpha=w_{1}\times{(r_{min})}_{overlap}+w_{2}\times{(r_{min})}_{others}$ (4)
$\beta=w_{1}\times{(r_{max})}_{overlap}+w_{2}\times{(r_{max})}_{others}$ (5)
@@ IV-B Quantization Method for Epitome
IV-C Architecture and Data Path of EPIM
@@
In practice, this means selecting the right inputs from the buffer and positioning them on the appropriate word lines of PIM crossbars. Also, voltages for those word lines connected to weights not part of this computation segment should be set to zero. Similarly, the outputs from the bit line must fit correctly within the output feature map. To efficiently manage this without adding to runtime, we’ve enhanced the base PIM accelerators with three index tables: the Input Feature Address Table (IFAT) and Input Feature Row Table (IFRT) for the input features, and the Output Feature Address Table (OFAT) for the outputs, as shown in Figure 2(b). The remaining PIM accelerator components remain consistent with existing work [10].
@@
Figure 3: The parameter size, latency, and energy consumption correspond to different layers (Layer 9, 41, 67) in ResNet50, with or without epitome.
V EPIM Design Optimization
V-A Motivation
@@
V-B Layer-Wise Epitome Design
@@
Figure 4: Comparison of the uniform-epitome solution with two optimization methods: output channel wrapping (marked as EPIM-Channel Wrapping) and layer-wise epitome design(marked as EPIM-Evo-Search), each used individually. EPIM-Opt refers to a combination of these two methods.
@@
$\small Reward=\frac{1}{\text{Latency}(E)}\times m\quad\text{or}\quad\frac{1}{\text{Energy}(E)}\times m,$ (6)
$\text{where}\quad m=\quad\begin{cases}0,&\text{if \#Crossbar}(E)>\text{Budget},\\ 1,&\text{if \#Crossbar}(E)\leq\text{Budget}.\end{cases}$ (7)
@@
V-C Output Channel Wrapping
@@
$\small W[x,:,:,:]=W[x+c,:,:,:],\quad x=0,\ldots,c\times(r-1)$ (8)
@@
$\small OFM[x,:,:,:]=OFM[x+c,:,:,:],\quad x=0,\ldots,c\times(r-1)$ (9)
@@
Table I: Experimental results of EPIM on ImageNet. CR stands for compression rate, XB stands for crossbar array, and mp stands for mixed-precision. The dimensions of an epitome are represented by a product notation. For example, “$1024\times 256$” represents that $c_{in}\times p\times q=1024$ and $c_{out}=256$
@@
EPIM-ResNet50-Latency-Opt W9A9 layer-wise 73.53 1080 12.15 49.2 16.4 93.4
EPIM-ResNet50-Energy-Opt W9A9 layer-wise 72.30 1048 12.52 50.6 15.6 93.2
@@
VI Experiments
VI-A Experimental Setup
For hardware performance evaluation, we develop a simulator based on MNSIM [9], a behavior-level simulator. We adapt the look-up table which is used for the latency and power parameters associated with basic hardware behaviors from MNSIM. The lookup table is modified to include the parameters of epitome and support channel wrapping. We use the 2-bit memristor cells in simulation. For the training of neural networks, we choose ResNet [11] as backbone algorithm. All the experiments are conducted on ImageNet [12]. Epitome-based neural networks and CNNs are both trained from scratch under full-precision. For quantization, we integrate the popular HAWQ [13] algorithm into our simulator. For comparison with pruning algorithms, we reproduced PIM-Prune, a pruning framework designed for neural networks on PIM accelerators[4].
VI-B Results and Analyses
The main experiment results of EPIM are summarized in Table I and Figure 4.
Table I provides a comparison between conventional NN algorithms and our EPIM solutions. We choose ResNet-50 and ResNet-101 to demonstrate the experimental results. For ResNet50, our method achieves up to a 30.65×30.65$\times$30.65 × compression rate of crossbars with a maximum accuracy loss of $4.78\%$. For ResNet-101, our method achieves up to a 31.22×31.$2$ compression rate of crossbars with a maximum accuracy loss of $3.79\%$. Combined with ultra-low bit quantization, EPIM reduces the inference energy by 23.01 $\times$ for ResNet-50 and 22.64$\times$ for ResNet-101. Our experiments show that epitomes, when integrated with quantization, are effectively deployable on PIM accelerators, which mainly support fixed-point computation. Additionally, our method outperforms PIM-Prune [4] in compression rate and accuracy under same quantization setting.
In Figure 4, we contrast manually designed uniform epitomes with epitomes produced by design optimization methods. Both layer-wise epitome design and output channel wrapping optimize the latency and energy of epitomes when applied individually. Layer-wise epitome design outperforms output channel wrapping in most settings. The EPIM-Opt method, combining output channel wrapping and layer-wise epitome design, outperforms all others. Compared to uniform epitomes, it offers similar compression with up to 3.07× speedup, 2.36× energy savings, and 7.13× reduced EDP. A comparison of ResNet50 under W9A9 quantization in Table I reveals that EPIM-ResNet50-Latency-Opt exhibits the lowest latency, while EPIM-ResNet50-Energy-Opt presents the lowest energy consumption, both without significant accuracy degradation.
VII Ablation Study
VII-A Detailed Experiments for Quantization
The detailed results of quantization are provided in Table II. We retrained six quantized models under two ultra-low-bit quantization scenarios. The results show that both of the proposed updates contribute to the increase in accuracy. For instance, for ResNet-50 under 3-bit quantization, the accuracy can be improved from 69.95% to 71.35% by adjusting scaling factor with crossbars. Moreover, by using the weighted sum scaling factor, the accuracy can be further improved to 71.59%.
Table II: Detailed Quantization Experiments Results.
@@ VII-A Detailed Experiments for Quantization
VII-B Detailed Comparison of Epitome and Pruning
In order to assess the epitome’s compatibility with pruning, we experiment with merging both concepts using basic element-wise pruning methods. Due to challenges in determining the crossbar compression rate with pruning, we compare parameter compression rates. Results are in Table III.
Table III: Accuracy and compression rate of epitome, epitome + element pruning with 50% pruning ratio, PIM-Prune[4].

### v3

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: 61st ACM/IEEE Design Automation Conference; June 23–27, 2024; San Francisco, CA, USAbooktitle: 61st ACM/IEEE Design Automation Conference (DAC ’24), June 23–27, 2024, San Francisco, CA, USAdoi: 10.1145/3649329.3657377isbn: 979-8-4007-0601-1/24/06conference: Design Automation Conference 2024; June 2024; San Francisco, CA, USAprice: 15.00isbn: 978-x-xxxx-xxxx-x/xx/xxdoi: 10.1145/nnnnnnn.nnnnnnn
1. Introduction
The emergence of Processing-In-Memory (PIM) accelerators offers a promising approach to boost the computational efficiency of neural network operations (Chi et al., 2016). Leveraging the capability to conduct computations directly within memory and eliminating the need for frequent data transfers between memory and computing units, PIM accelerators offer a substantial improvement in energy efficiency. Remarkably, these accelerators surpass the performance of GPU and CMOS ASIC solutions by more than 100-fold (Shafiee et al., 2016; Chi et al., 2016).
Despite the advances in current PIM accelerators, they still face certain challenges. One challenge arises from the higher cost of writing to emerging memory compared to reading from it. Due to this limitation, PIM accelerators typically require loading all neural network weights onto memristor crossbars prior to conducting computations. However, modern large-scale neural network models have seen a significant increase in parameter sizes. Even for the existing PIM accelerators with maximum memristors, there still exists a significant challenge in deploying a standard backbone model like ResNet-50(Wan et al., 2022). As such, the deployment of advanced Convolutional Neural Networks (CNN) models on PIM is hindered.
To address this challenge, researchers have proposed methods that fall into two main categories. Various methods, including pruning and quantization, aim to represent neural operator weights using a reduced number of parameters(Sun et al., 2020; Chu et al., 2020). Conversely, approaches such as neural architecture search concentrates on identifying optimal combinations of neural operators(Jiang et al., 2020; Sun et al., 2022). Despite the advancements, the majority of these studies have focused on adapting existing convolution operators. The crucial endeavor of developing compact and high-performance operators specifically tailored for PIM accelerators has largely been overlooked.
@@ 1. Introduction
To craft such neural operator, we began by reviewing existing literature. Prior research highlights the potential of leveraging overlaps in convolution kernels, introducing a compact neural operator termed “epitome”(Zhou et al., 2019). This operator reconstructs convolution kernels from a smaller epitome tensor. A more compact neural network can be crafted by replacing the its convolutions with epitomes. We identified that, with software and hardware adjustments, epitome could be adapted to PIM accelerators. Specifically, from the hardware side, epitome only introduces minor modifications to the data path. Hence, we adopted epitome as our foundation. By integrating epitome with our proposed methods, we achieved notable compression rates while maintaining a high accuracy.
@@
2. Related work
2.1. Process-In-Memory Architecture for CNN Acceleration
Commonly, PIM architecture consists of multiple memristor crossbars. The weight matrices of CNN are mapped as the conductance of the memristor crossbar cells. However, the writing latency of the memristor crossbar cell is multiple times larger than the reading latency(Chi et al., 2016). Before inference, the weight matrices of all CNN layers need to be mapped onto memristor crossbars. Such restrictions stimulate the need to support model compression algorithms on PIM architecture.
Figure 1. Epitome’s Reconstruction into Convolution
2.2. Epitome
@@
(1) $\small E_{s}=E_{i}[p:p+w,q:q+h,c_{in}:c_{in}+\beta_{1},c_{out}:c_{out}+\beta_{2}],$
where $(p,q,c_{in},c_{out})$ denote the starting index of the sub-tensor in the epitome and $(w,h,\beta_{1},\beta_{2})$ denotes the length of the sub-tensor along each dimension. The sampling process is repeated until the fetched small pieces of weight tensors can be concatenated to match the dimensions of the convolution. Figure 1 is a visual illustration of such process. Unlike convolutions, whose shapes are determined by the input and output channels, epitomes have greater flexibility in shape. They can be of four-dimensional tensors with any size, as long as their sampled patches can reconstruct the convolutions(Zhou et al., 2019).
2.3. Quantization
Quantization methods reduce model size by using low bitwidth for weights and activations in neural networks(Dong et al., 2020). In quantization, the quantizer maps weights and activations into integers with a real-valued scaling factor $S$ and an integer zero point $Z$ as follows:
(2) $\small Q(r)=\text{Int}\big{(}{r}/{S}\big{)}-Z$
@@ 2.3. Quantization
(3) $\small S=\frac{\beta-\alpha}{2^{k}-1}$
@@
3. Framework Overview
Figure 2. Illustration of EPIM Framework.
@@ 3. Framework Overview
4. EPIM Design Method
This section introduces our method EPIM. Section 4.1 shows the adjustments of epitome on PIM accelerators. Section 4.2 presents details of our quantization method. Section 4.3 describes the specifications of our PIM accelerator.
4.1. Mapping and Adjusting Epitomes for PIM Accelerators
Just like convolutions, epitomes are four-dimensional tensors that can be mapped onto memristors in a similar manner. The distinction between epitomes and convolutions arises during inference: while all word lines and bit lines corresponding to a convolution’s weight are activated all together, an epitome causes a memristor crossbar to be activated multiple times. Each activation only engages the word lines and bit lines associated with small patches sampled from the epitome. In our implementation, we use the same mapping strategy as (Zhu et al., 2020).
Motivated by the size flexibility of the epitomes, we can adjust their shapes to better utilize memristors. Specifically, we aim for $c_{out}$ and $c_{in}\times p\times q$ to align as integral multiples of the crossbar size. Following the mapping approach from (Zhu et al., 2020), where $c_{in}$, $p$ and $q$ dimensions correspond to word lines, while $c_{out}$ dimension corresponds to bit lines on memristor crossbars. By setting to these dimensions to desired values, we ensure epitomes make full use of memristor crossbars.
4.2. Quantization Method for Epitome
When directly quantizing the epitome with ultra-low precision, we observed a notable decrease in accuracy. Table 2 indicates that while the 5-bit quantized weights for ResNet-50’s epitome achieve 73.59% accuracy, this drops to 69.95% for 3-bit weights. To enhance accuracy under such low-bit quantization, we suggest two adjustments: Firstly, given the parallel computation between PIM accelerator crossbars, we allocate a scaling factor to each crossbar. Secondly, we observe that the epitome reconstructs convolution by sampling overlapping small patches, with certain patches from the epitome being repeated more often than others in the reconstructed convolution. When quantizing the epitome, it’s logical to assign different importance levels to different parts based on their repetition frequency. As depicted in Figure 2(c), we notice that the center parts (green) of epitomes are usually repeated more frequently than the other parts (blue). We calculate the weighted sum of the epitome’s maximum and minimum values in the center parts and other parts as quantization range $\alpha$ and $\beta$, based on these repetitions. The detailed calculation process is as follows:
(4) $\small\alpha=w_{1}\times{(r_{min})}_{overlap}+w_{2}\times{(r_{min})}_{others}$
(5) $\beta=w_{1}\times{(r_{max})}_{overlap}+w_{2}\times{(r_{max})}_{others}$
@@ 4.2. Quantization Method for Epitome
4.3. Architecture and Data Path of EPIM
@@
In practice, this means selecting the right inputs from the buffer and positioning them on the appropriate word lines of PIM crossbars. Also, voltages for those word lines connected to weights not part of this computation segment should be set to zero. Similarly, the outputs from the bit line must fit correctly within the output feature map. To efficiently manage this without adding to runtime, we’ve enhanced the base PIM accelerators with three index tables: the Input Feature Address Table (IFAT) and Input Feature Row Table (IFRT) for the input features, and the Output Feature Address Table (OFAT) for the outputs, as shown in Figure 2(b). The remaining PIM accelerator components remain consistent with existing work (Zhu et al., 2019).
@@
5. EPIM Design Optimization
Figure 3. The parameter size, latency, and energy consumption correspond to different layers (Layer 9, 41, 67) in ResNet50, with or without epitome.
5.1. Motivation
@@
5.2. Layer-Wise Epitome Design
@@
Figure 4. Comparison of the uniform-epitome solution with two optimization methods: EPIM-Channel Wrapping and EPIM-Evo-Search, each used individually. EPIM-Opt refers to a combination of these two methods.
@@
(6) $\small Reward=\frac{1}{\text{Latency}(E)}\times m\quad\text{or}\quad\frac{1}{\text{Energy}(E)}\times m,$
(7) $\text{where}\quad m=\quad\begin{cases}0,&\text{if \#Crossbar}(E)>\text{Budget},\\ 1,&\text{if \#Crossbar}(E)\leq\text{Budget}.\end{cases}$
@@
5.3. Output Channel Wrapping
@@
(8) $\small W[x,:,:,:]=W[x+c,:,:,:],\quad x=0,\ldots,c\times(r-1)$
@@
(9) $\small OFM[x,:,:,:]=OFM[x+c,:,:,:],\quad x=0,\ldots,c\times(r-1)$
@@
Table 1. Experimental results of EPIM on ImageNet. CR stands for compression rate, XB stands for crossbar array, and mp stands for mixed-precision. The dimensions of an epitome are represented by a product notation. For example, “$1024\times 256$” represents that $c_{in}\times p\times q=1024$ and $c_{out}=256$
@@
EPIM-ResNet50-Latency-Opt W9A9 layer-wise 73.60 1080 12.15 49.2 16.4 93.4
EPIM-ResNet50-Energy-Opt W9A9 layer-wise 73.15 1048 12.52 50.6 15.6 93.2
@@
6. Experiments
6.1. Experimental Setup
The development of the simulator is based on MNSIM (Zhu et al., 2020), a highly accurate, behavior-level simulator. Following the practice in MNSIM, we maintain a look-up table for the storage of the latency and power parameters associated with basic hardware behaviors. The lookup table is modified to include the parameters of epitome. We use the well-explored 2-bit memristor cells in simulation. For quantization, we integrate the popular HAWQ (Dong et al., 2019) algorithm into our simulator. All the experiments are conducted on the large-scale benchmark ImageNet (Deng et al., 2009). For comparison purposes, we reproduced PIM-Prune, a pruning framework designed for neural networks on PIM accelerators(Chu et al., 2020).
6.2. Results and Analyses
The main experiment results of EPIM are summarized in Table 1 and Figure 4.
Table 1 provides a comparison between conventional NN algorithms and our EPIM solutions. We choose ResNet-50 and ResNet-101 to demonstrate the experimental results. For ResNet50, our method achieves up to a 30.65×30.65$\times$30.65 × compression rate of crossbars with a maximum accuracy loss of $4.78\%$. For ResNet-101, our method achieves up to a 31.22×31.$2$ compression rate of crossbars with a maximum accuracy loss of $3.79\%$. More importantly, combined with ultra-low bit quantization, EPIM reduces the inference energy by 23.01 $\times$ for ResNet-50 and 22.64$\times$ for ResNet-101. Our experiments show that EPIM can serve as a competitive solution to achieve efficient PIM deployment of large CNNs.
In Figure 4, we contrast manually designed uniform epitomes with epitomes produced by design optimization methods. The EPIM-Opt method, combining Channel Wrapping and Evo-Search, outperforms others. Compared to uniform epitomes, it offers similar compression with up to 3.07× speedup, 2.36× energy savings, and 7.13× reduced EDP. A comparison of ResNet50 under W9A9 quantization in Table 1 reveals that EPIM-ResNet50-Latency-Opt exhibits the lowest latency, while EPIM-ResNet50-Energy-Opt presents the lowest energy consumption, both without significant accuracy degradation.
7. Ablation Study
7.1. Detailed Experiments for Quantization
The detailed results of quantization are provided in Table 2. We retrained six quantized models under two ultra-low-bit quantization scenarios. The results show that both of the proposed updates contribute to the increase in accuracy. For instance, for ResNet-50 under 3-bit quantization, the accuracy can be improved from 69.95% to 71.35% by adjusting scaling factor with crossbars. Moreover, by using the weighted sum scaling factor, the accuracy can be further improved to 71.59%.
Table 2. Detailed Quantization Experiments Results.
@@ 7.1. Detailed Experiments for Quantization
7.2. Detailed Comparison of Epitome and Pruning
In order to assess the epitome’s compatibility with pruning, we experiment with merging both concepts using basic element-wise pruning methods. Due to challenges in determining the crossbar compression rate with pruning, we compare parameter compression rates. Results are in Table 3.
Table 3. Accuracy and compression rate of epitome, epitome + element pruning with 50% pruning ratio, PIM-Prune(Chu et al., 2020).